OpenStudy (anonymous):

Need help to find the volume of a sphere. The formula for volume of a sphere is : V=4/3 *pi*r^3 ( i think )

OpenStudy (anonymous):

Find the volume of the sphere to the nearest whole number. Use p = 3.14.

OpenStudy (anonymous):

I used the formula and this is what i got : v=4/3*pi*5^3 = 523.5. Is this the correct answer ?

OpenStudy (agent47):

yes.. but u need to round it to the nearest whole number, so: 523

OpenStudy (anonymous):

Round to the nearest whole number gives 524

OpenStudy (anonymous):

i got 1500 because you 5^3=5*5*5=125

OpenStudy (anonymous):

v=4/3*3.14*125 v=?

OpenStudy (agent47):

well the answer was 523.3333, not 523.5, so rounding gave me 523

OpenStudy (anonymous):

oh.. okay i see what i did and yes you are correct

OpenStudy (anonymous):

Thank you all for replying.